OpenStudy (anonymous):

Solve the Equation n-19=-16

OpenStudy (uri):

n=-16+19

OpenStudy (anonymous):

i got n=3

OpenStudy (anonymous):

how did you both get those answers???

OpenStudy (anonymous):

Add \(19\) both the sides: \[n - 19 + 19 = -16 + 19\]

Nnesha (nnesha):

to cancel the -19 on left side to get n by itself add 19 both side

OpenStudy (anonymous):

\[n \cancel{- 19} + \cancel{19} = -16 + 19 \implies \color{green}{n = 3}\] Getting?

Nnesha (nnesha):